OpenStudy (anonymous):

calculate integrate(x^(x))dx,x=0..1)

OpenStudy (anonymous):

Integrate(ln(x)*ln(1-x))dx,x=0 to 1

OpenStudy (unklerhaukus):

\[\int\limits_0^1\ln(x)\ln(1-x)\text dx\]